Two sides of a triangle are equal. Each of the two sides is twice the length of the third side. The perimeter is 90 meters. Find the lenght of each side.
Answer by ankor@dixie-net.com(12692)

Two sides of a triangle are equal. Each of the two sides is twice the length of the third side. The perimeter is 90 meters. Find the length of each side.
:
Let the 3rd side = x
Then the equal sides = 2x
:
x + 2x + 2x = 90
5x = 90
x = 90/5 = 18 is the length of the 3rd side.
:
Equal sides: 2(18) = 36
:
Check solution: 18 + 36 + 36 = 90